# Leveraging Large Language Models for Relevance Judgments in Legal Case Retrieval

### v1

@@ 3.3. Other Statements
Apart from that, the data of Chinese legal documents includes more than just the ”basic case information”, but we only utilize this part for relevance annotation. There are two main reasons for this:
@@
  • Consistency with Dataset Standards: By focusing on case facts, our approach aligns with the Lecard dataset’s method of annotation, which is based on key facts (MF and LF) within case details. This ensures that our method aligns with the dataset’s standards for relevance annotation.
3.4. Application of annotated data
An application of our method is data augmentation: to construct a synthetic dataset from unlabeled legal cases for model training. Take LeCaRD as an example, where there are a total of 10,700 candidate cases. By randomly selecting two candidate cases each time for relevance annotation, theoretically, we can generate over 50 million pairs of data. This could significantly scale the annotated dataset. We tried this on a small scale: We first randomly generated 200k candidate case pairs. Our goal is to produce more high-quality positive examples. Considering that the case pairs randomly extracted have a high probability of being irrelevant, the 200k case pairs were pre-sorted by a BERT model fine-tuned on the original dataset (see the Experiment section for implementation details) and we performed relevance annotation on the top 50k data. After obtaining these annotations, we constructed two augmented datasets intended for fine-tuning the retrieval models: the first, a 20k dataset with a distribution that closely mirrors the label distribution of the original dataset; and the second, a 40k dataset selected randomly from the 50k annotations.
@@ 4. Experiment
In this section, we conduct various empirical experiments to answer the following research questions (RQs):
RQ 1 How reliable are the relevance judgments obtained with the proposed workflow?
RQ 2 How do the LLM-based relevance judgments align with the judgments from human annotators?
RQ 3How can we leverage the LLM-based relevance judgments to augment existing legal case retrieval models?
@@ 4.1. Datasets and Evaluation Metrics
Table 1. The detail of the dataset.
Description Value
DATASETS LeCaRD
Language Chinese
# Train query case -
# Train candidate cases/query -
# Test query case 107
# Test candidate cases/query 100
Avg. length per case document 8,275
# Avg. relevant cases per query 10.33
Table 2. Statistics for the golden labels. In the dataset, each query has 100 candidates, with only the top 30 candidates having a golden label of relevance, while the remaining 70 candidates are irrelevant.
Label Mean (Percentage) Median Max Min
0 3.94 (13.15%) 2.0 28 0
1 5.96 (19.88%) 4.0 29 0
2 9.87 (32.90%) 9.0 30 0
3 10.39 (34.64%) 8.0 30 1
@@
We carried out our experiments using the LeCaRD (Chinese Legal Case Retrieval Dataset) (Ma et al., 2021). This dataset comprises more than 43,000 candidate cases and 107 query cases. All the queries and results are sourced from criminal cases made public by the Supreme People’s Court of China. Notably, while only the fact paragraph is present in the LeCaRD queries, the candidate documents encompass the full case. For every query, there is a set of 100 corresponding candidate cases. In legal case retrieval, relevance is categorized with multi-level labels (0 to 3). For a pair of cases: 0 indicates no relevance, 1 indicates that the Material Facts are relevant but the Legal Facts are not, 2 indicates that the Legal Facts are relevant but the Material Facts are not, and 3 indicates relevance to both kinds of facts. The detail of this dataset is shown in Figure 1 and Figure 2.
@@
To the best of our knowledge, our method is the first automatic Chinese legal case annotation approach, therefore, to evaluate our method, we compare the annotation consistency with the human-annotated labels by Cohen’s Kappa. In addition, we choose normalized discounted cumulative gain (NDCG) for evaluating the performance of legal case retrievers before and after training on the augmented data, to indirectly illustrate the effectiveness of our method.
4.2. Evaluation of Annotations.
Separately based on two aspects MF and LF, we assess the reliability of the proposed annotation method by calculating Cohen’s Kappa for multiple annotations conducted at a certain temperature. Following that, we evaluate the automated annotation consistency to the human labels by calculating Kappa values with the golden labels. Finally, we draw heatmaps to visualize the consistency between the proposed method and human labeling, to demonstrate the validity.
For the reliability shown in Fiture 9, as the temperature increases, there is a moderate change in the Kappa values among our multiple relevance annotations, indicating good reliability and that our carefully designed expert instruction can effectively orient the output of LLM. For the validity shown in Fiture 10, it can be observed that the consistency between Legal Facts and human labels is highly significant, while Material Facts exhibit relative significance when compared to human labels. One primary reason for this is that each case involves one or more defendants; when comparing the facts related to the defendants between two cases, there are one-to-many and many-to-many comparative relationships, which could involve more subjective judgment. During our research, we do identify a significant degree of subjective judgment in the human annotation of Material Facts under such comparisons. This is attributed to the fact that human labels do not originate from a single individual, and there is inevitably some bias among different individuals’ judgments. Moreover, this bias is more likely to manifest in Material Fact, due to the arbitrary forms of writing it. However, according to the labeling standard, the relevance of legal facts (score=2) is more critical than that of material facts (score=1). Therefore, the quality of annotations for Legal Facts still predominantly influences the final annotation results. This indicates that our proposed automatic annotation method can closely approximate human labeling. Interestingly, in all temperature settings, the avg. kappa values remain relatively stable. This illustrates the reliability of aligning with human labels again.
The heatmap in Figure 11 and Figure 12 shows the high consistency of our automated annotation approach with human labeling, especially in cases of non-relevance and full relevance. This further demonstrates the validity of the annotation results.
Figure 9. The Cohen’s Kappa between multiple times of MF&LF annotation results over different temperature settings: The x-axis represents the temperature, while the y-axis indicates Cohen’s Kappa value. The Kappa value of multiple generated results indicates the reliability of the method’s output.
Figure 10. The Cohen’s Kappa between the MF&LF and the golden labels over different temperature settings: The x-axis represents the temperature, while the y-axis indicates Cohen’s Kappa value. ’Avg.’ represents the average of Kappa values between multiple annotation results and the golden labels, measuring the consistency of our annotations with golden labels, which indicates the validity of our approach.
4.3. Baselines for Data Augmentation Experiments
To further validate the annotation quality, we examine the improvements achieved when using annotated synthetic data for training different retrievers in LCR tasks. This will further explore the effectiveness of this automatic annotation method, which will be detailed in subsequent chapters. Here, we select only two basic deep models but not more complicated models like (Li et al., 2023). Since our goal is to compare the relative differences in model performance before and after using the proposed annotation for data augmentation, to validate whether downstream models can be effectively augmented, rather than pursuing a state-of-the-art model for the LCR task.
  • BM25 (Robertson and Walker, 1994) is a simple but effective baseline for LCR.
  • BERT (Devlin et al., 2018) is built on the Transformer architecture, which employs self-attention mechanisms to weigh input tokens differently, allowing for a more nuanced representation of context.
  • Longformer (Beltagy et al., 2020) is an adaptation of the Transformer architecture designed to process exceptionally long sequences efficiently. It achieves this by combining local attention with a limited number of global attention mechanisms, reducing the computational complexity from quadratic to linear for most tokens. This design enables direct processing of tasks involving extensive documents without the need for chunking or hierarchical methods.
    Figure 11. The heatmap of the annotated four-level relevance vs golden labels.
    Figure 12. The heatmap of the annotated MF&LF relevance vs golden labels.
@@ 4.4. Implementation Details
All experiments are conducted on the LeCaRD dataset. The size of the demonstration set we built is 40 groups. Each group consists of a query case and four candidate cases with different relevance levels, which are evaluated in both FE and FA. For the annotation evaluation, we compare the annotation of the proposed method with the golden labels of the top 30 candidate cases in every query case. In the heatmap experiment, we randomly sample 100 human-labeled pairs for each relevance, from 0 to 3, 400 in total. For the data augmentation experiments, we split LeCaRD in a ratio of 8:1:1 for training, validation, and testing. The augmented data pairs are sampled from the candidate case corpus from the training set. We use PyTorch (Paszke et al., 2019) to implement our method. We use Nvidia 3090 with 24G GPU memory. We use the Chinese tokenizer Jieba  [^1] https://github.com/fxsjy/jieba in data pre-processing and use AdamW (Loshchilov and Hutter, 2019) optimizer in fine-tuning procedure. All hyper-parameters are tuned based on the performance of the validation set. Specifically, when using the BERT baseline model, We adopted a segmentation method similar to BERT-PLI, splitting long texts into chunks of length 256 with overlapping window sizes of 128. The [CLS] token serves as the embedding for each block, and during relevance computation, we cross-calculate the dot product of the query and candidate block vectors. Furthermore, the LeCaRD is in Chinese, so we use the Chinese version of Longformer, longformer-chinese. For large language models, we opted for the robust GPT-3.5-turbo  [^2] https://platform.openai.com/docs/models/gpt-3-5.
4.5. Results of Data Augmentation Experiments
Table 3. The performance of different baselines after pre-training on generative data of various sizes in the case retrieval task. The numbers following the models represent the size of the generated dataset used for pre-training.
MODEL&DATA NDCG@30
BM25 0.8532
BERT 0.8816
BERT+2k 0.8979
BERT+20k 0.9305
BERT+40k 0.9156
Longformer 0.9019
Longformer+20k 0.9458
@@ 4.5. Results of Data Augmentation Experiments
Table 4. The impact of different components for annotation quality: Original indicates the proposed approach. w/o ADM indicates the method of selecting examples from the demo library using random sampling; w/o FE means letting the LLM extract the two types of facts directly in the prompting workflow without prompt guidance and demo examples; w/o FA indicates that after extracting the two types of facts from each case, the similarity is directly labeled without the guidance of positive and negative demos in the original FA progress.
Annotation methods MF Kappa LF Kappa
Original 0.34 0.53
w/o ADM 0.23 0.54
w/o FE 0.09 0.21
w/o FA 0.19 0.50
w/o FE&FA 0.13 0.17
Table 5. The impact of different components for augmentation effectiveness.
Model NDCG@30
BERT2K 0.8979
BERT2K w/o ADM 0.8874
BERT2K w/o FE 0.8548
BERT2K w/o FA 0.8812
BERT2K w/o FA&FA 0.8501
Table 3 illustrates the performance of several baseline models on the task of legal case retrieval. We compared the performance changes of Bert and Longformer before and after fine-tuning on synthetic data, with BM25 retrieval results only used as a performance reference here. According to the results, we have some observations as follows: We pre-trained the BERT-based case retrieval model on three different scales of synthetic datasets. The 2k and the 20k datasets have a similar positive-to-negative ratio based on the distribution of the training set. The 40k dataset is randomly collected from all generated data. It can be observed that both BERT and Longformer, when trained on the 2k synthetic dataset of the same scale as the training set, already achieve significant improvements. The performance further significantly improves when trained on the larger 20k synthetic dataset. This demonstrates the data augmented by our method has effectively simulated real data. Intriguingly, even with twice the data expansion, the performance on the 40k dataset was notably worse than on the 20k dataset. There might be two explanations for this: A training dataset closer to the testing dataset in distribution benefits the model more significantly. Random sampling without selection or filtering retains fewer positives and more unnecessary negatives. An excessive number of negative examples will not necessarily bring more value to the model.
4.6. Ablation Study
@@
In this section, we aim to explore the impact of each stage in our proposed data annotation workflow on the final quality of the generated data. The ablation study is conducted for both annotation and augmentation.
@@
According to Table LABEL:aba, removing ADM impacts the annotation of MF notably. This indicates that the extraction and relevance assessment of Material Facts relies on more pertinent expert guidance. This may be due to the variable nature of Material Facts. Meanwhile, Legal Facts remain unaffected due to their normative formats. This also highlights that the understanding of Legal Facts in the case benefits more from the FE and FA modules. When FE is removed, the Kappa scores for both types of facts significantly decrease, and the effect is similar when both FE and FA are removed. This demonstrates that the fact extraction step we designed is the most crucial part of the entire annotation process. Based on this observation, it can be inferred that LLM faces difficulty in accurately identifying key legal elements. Removing FA has little impact on LF annotation, but it does negatively affect MF. This indirectly confirms that extracting MF from case pairs and assessing their relevance is a challenging aspect of judging legal case relevance, and our method addresses this challenge through a series of carefully designed processes.
For the augmentation ablation shown in Table LABEL:Table_4, when not using ADM, the randomly selected demo might differ significantly from the current query case in content. This weakens the guiding role of expert knowledge for the LLM during information extraction and annotation, leading to a noticeable impact on the performance of the downstream task model. Also, by eliminating the expert knowledge guidance in the FA phase, the performance of the downstream model is similarly weakened. We speculate that the criteria for similarity judgment in the last step are clearer with expert guidance. The performance of the downstream model is greatly affected when FE is removed, even worse than when no synthetic data is applied. We believe this is because, when identifying the two Legal Facts that have a decisive impact on case similarity, the LLM struggles to independently recognize and judge key legal elements and their interconnections. Our few-prompt method inspires the LLM to mimic experts.
@@
Every stage within the prompt workflow contributes significantly to enhancing the quality of domain data annotation. Both Adaptive Demo-Matching (ADM) and Few-shot Annotation (FA) moderately influence the quality of annotations, with the impact of ADM being more pronounced. ADM validates that relevant demonstrations indeed provide LLMs with a more robust basis for reasoning. On the other hand, FA introduces clearer constraint conditions for relevance annotations. Notably, the most pivotal process is FE, involving the most specialized knowledge Furthermore. It also suggests that our FE method can effectively leverage minimal expert instructions to inspire and guide LLMs in capturing subtle but vital nuances within professional data.

### v2

@@ 3.3. Other Statements
Apart from that, the data of Chinese legal documents includes more than just the ”basic case information”, but we only utilize this part for relevance judgments. There are two main reasons for this:
@@
  • Consistency with Dataset Standards: By focusing on case facts, our approach aligns with the Lecard dataset’s method of judgments, which is based on key facts (MF and LF) within case details. This ensures that our method aligns with the dataset’s standards for relevance annotation.
3.4. Utility of the Relevance Judgments
To further analyze the utility of our relevance judgments, we use the data annotated with relevance judgments to perform data augmentation for downstream models: to construct a synthetic dataset from unlabeled legal cases for model training. Take LeCaRD as an example, where there are a total of 10,700 candidate cases. By randomly selecting two candidate cases each time for relevance annotation, theoretically, we can generate over 50 million pairs of data. This could significantly scale the annotated dataset. We tried this on a small scale: We first randomly generated 200k candidate case pairs. Our goal is to produce more high-quality positive examples. Considering that the case pairs randomly extracted have a high probability of being irrelevant, the 200k case pairs were pre-sorted by a BERT model fine-tuned on the original dataset (see the Experiment section for implementation details) and we performed relevance annotation on the top 50k data. After obtaining these annotations, we first constructed two synthetic label-only datasets intended for fine-tuning the retrieval models: the first, a 2k and a 20k dataset with distributions that closely mirror the label distribution of the original dataset; and the second, a 40k dataset selected randomly from the 50k annotations. Then we use the 2k synthetic dataset, along with the LLM analysis process saved during its generation, to fine-tune open-source small LLMs, enhancing their capability of legal relevance judgments.
@@ 4. Experiment
In this section, we conduct various empirical experiments to answer the following research questions:
RQ 1: Does the workflow reliably produce consistent and reasonable judgments?
RQ 2: How do the LLM-based relevance judgments align with the judgments from human annotators?
RQ 3: How can LLM-based relevance judgments be utilized to enhance existing legal case retrieval models and to train open-source large models for assessing case similarity?
RQ 4: How much does each component of our workflow contribute to the overall judgment accuracy?
Briefly, RQ1 focuses on internal workflow consistency and reliability, while RQ2 focuses on external alignment with human standards. Both questions are important and complementary, as they address different aspects of the workflow’s performance.
@@ 4.1. Datasets and Evaluation Metrics
We carried out our experiments using the LeCaRD (Chinese Legal Case Retrieval Dataset) (Ma et al., 2021). This dataset comprises more than 43,000 candidate cases and 107 query cases. All the queries and results are sourced from criminal cases made public by the Supreme People’s Court of China. Notably, while only the fact paragraph is present in the LeCaRD queries, the candidate documents encompass the full case. For every query, there is a set of 100 corresponding candidate cases. In LeCaRD, relevance is categorized with multi-level labels (0 to 3). For a pair of cases: 0 indicates no relevance, 1 indicates that the Material Facts are relevant but the Legal Facts are not, 2 indicates that the Legal Facts are relevant but the Material Facts are not, and 3 indicates relevance to both kinds of facts.
@@
To the best of our knowledge, our method is the first for automatic Chinese legal case relevance judgments. To evaluate our method, we compare the judgment consistency with the human-annotated labels by Cohen’s Kappa. In addition, we choose normalized discounted cumulative gain (NDCG) for evaluating the performance of legal case retrievers before and after training on the augmented data, to indirectly illustrate the effectiveness of our method.
@@ 4.2. Implementation Details
According to the crime classification defined in the chapters of the Criminal Law’s specific provisions, the demonstration set in ADM is constructed based on 40 query cases, which cover the main types of cases present in the LeCaRd dataset. For each query case, four associated cases with different gold-standard relevance labels are provided, along with FE and FA instructions corresponding to two factual aspects: MF and LF.
For the relevance evaluation, we compare the judgments of the proposed method with the golden labels of the top 30 candidate cases in every query case. For the LLM, we opted for GPT-3.5-0613 with temperature 0.4. In the heatmap experiment, we randomly sample 400 human-labeled pairs for all relevance levels, from 0 to 3.
@@
For the data augmentation experiments, we split LeCaRD in a ratio of 8:1:1 for training, validation, and testing. The augmented data pairs are sampled from the candidate case corpus from the training set. We use the Chinese tokenizer Jieba  [^1] https://github.com/fxsjy/jieba in data pre-processing and use AdamW (Loshchilov and Hutter, 2019) optimizer in fine-tuning retrieval models. All hyper-parameters are tuned based on the performance of the validation set. Specifically, when using the BERT  [^2] https://huggingface.co/google-bert/bert-base-chinese baseline model, we adopted a segmentation method similar to BERT-PLI, splitting long texts into chunks of length 256 with overlapping window sizes of 128. The [CLS] token serves as the embedding for each block, and during relevance computation, we cross-calculate the dot product of the query and candidate block vectors. Furthermore, the LeCaRD is in Chinese, so we use the Chinese version of Longformer, longformer-chinese. For the LLMs, we set the temperature to 0.5. When fine-tuning open-source LLMs using 2k synthetic data of GPT-3.5 with its analysis process, we applied LoRA with a batch size of 256, a rank of 32, a learning rate of 5e-4, and a context window of 2048.
4.3. Evaluation of the Relevance Judgments (RQ1 & 2)
Separately based on two aspects, Material Facts (MF) and Legal Facts (LF), we assess the reliability of the proposed method by calculating Cohen’s Kappa for multiple judgments conducted at a certain temperature in Figure 9. To demonstrate the validity, we evaluate the automated judgment consistency by calculating Kappa values with the golden labels in Figure 10, and then draw heatmaps that visualize the consistency between the proposed method and human labeling.
For the reliability test shown in Figure 9, as the temperature increases, there is a moderate change in the Kappa values among our multiple relevance judgments, indicating good reliability and that our carefully designed expert instruction can effectively orient the output of LLM. For the validity test shown in Figure 10, it can be observed that the consistency between Legal Facts and human labels is highly significant, while Material Facts exhibit relative significance when compared to human labels. One primary reason for this is that each case involves one or more defendants; when comparing the facts related to the defendants between two cases, there are one-to-many and many-to-many comparative relationships, which could involve more subjective judgment. During our research, we do identify a significant degree of subjective judgment in the human judgments of Material Facts under such comparisons. This is attributed to the fact that human labels do not originate from a single individual, and there is inevitably some bias among different individuals’ judgments. Moreover, this bias is more likely to manifest in Material Fact, due to the arbitrary forms of writing it. However, according to the labeling standard, the relevance of legal facts (score=2) is more critical than that of material facts (score=1). Therefore, the quality of judgments for Legal Facts still predominantly influences the final judgment results. This indicates that our proposed automatic relevance judgment can closely approximate human labeling. Interestingly, in all temperature settings, the Avg. kappa values remain relatively stable. This illustrates the reliability of aligning with human labels again.
The heatmaps in Figure 11 and Figure 12 show the high consistency of our automated relevance judgments with human labeling, especially in cases of non-relevance and full relevance. This further demonstrates the validity of the judgment results.
Figure 9. The Cohen’s Kappa between multiple times of MF&LF relevance judgment over different temperature settings.
Figure 10. The Cohen’s Kappa between the MF&LF and the golden labels over different temperature settings: Avg. represents the average of Kappa values between multiple relevance judgments and the golden labels
4.4. Data Augmentation Experiments (RQ3)
To explore the annotation capabilities of different LLMs and whether the proposed workflow in this paper can further enhance the annotation capabilities of open-source light-weighted LLMs through fine-tuning. We first fine-tune various small LLMs using generated 2k synthetic data with analysis processes annotated by GPT-3.5. Then, using the proposed workflow with GPT-3.5 replaced by smaller LLMs, we compare their consistency with human-annotated golden labels on the test set. To further validate the utility of our automatic relevance judgment, we analyze the improvements achieved by training different retrieval models on annotated synthetic data in LCR tasks.
4.4.1. Baselines
For relevance judgment capabilities of different LLMs, we opt for GPT-3.5-0614 (referred to as GPT-3.5), Llama-2-7B-chat-hf(Touvron et al., 2023b) (referred to as Llama-2) and Qwen-2-7B-instruct(Bai et al., 2023) (referred to as Qwen-2).
In We select two basic deep models (BERT (Devlin et al., 2018), Longformer (Beltagy et al., 2020)) and an advanced SOTA method SAILER (Li et al., 2023b) as our baselines. Note that our goal is to evaluate the effectiveness of automatic relevance judgments. The focus is on validating whether the relative performance differences in downstream models before and after applying the augmented data, rather than achieving state-of-the-art results on the LCR task.
Figure 11. The heatmap of the four-level relevance judgments vs golden labels.
Figure 12. The heatmap of the relevance judgments MF&LF relevance vs golden labels.
4.4.2. Data Augmentation Results for Fine-tuning small LLMs
As shown in Table 1, using our proposed workflow and a small amount of synthetic data generated by the powerful general LLM GPT-3.5, the capability for legal relevance judgment of two 7B LLMs can be significantly improved, reaching or even surpassing that of GPT-3.5.
Table 1. Relevance consistency (Cohen’s Kappa) with golden labels in the test set, where * denotes models fine-tuned with synthetic data.
GPT-3.5 Llama-2 Qwen-2 Llama-2* Qwen-2*
LF 0.67 0.41 0.44 0.63 0.69
MF 0.62 0.25 0.29 0.55 0.60
4.4.3. Data Augmentation Results for Fine-tuning Retrieval Models
Table 2. NDCG@30 of different baselines after pre-training on label-only synthetic data of various sizes, generated by either the original or fine-tuned LLMs. G, L, and Q denote GPT-3.5, Llama-2, and Qwen-2, respectively, while * indicates fine-tuned LLMs with explanations from GPT-3.5. Bold numbers indicate the best performance in each row, while underlined numbers represent the second best. The 2k and 20k datasets have a similar positive-negative ratio to the training set, while the 40k dataset was randomly collected from all generated data.
Baseline Original G L Q L* Q*
BERT 0.8816 - - - - -
Longformer 0.9019 - - - - -
SAILER 0.9223 - - - - -
BERT+2k - 0.8979 0.8794 0.8851 0.8895 0.8967
BERT+20k - 0.9305 0.9043 0.9156 0.9241 0.9330
BERT+40k - 0.9156 0.8962 0.9007 0.9087 0.9125
Longformer+20k - 0.9458 0.9129 0.9283 0.9374 0.9480
SAILER+20k - 0.9606 - - 0.9479 0.9623
@@ 4.4.3. Data Augmentation Results for Fine-tuning Retrieval Models
Table 3. The impact of different components for relevance quality: Original indicates the proposed approach. w/o ADM indicates the method of selecting examples from the demo library using random sampling; w/o FE means letting the LLM extract the two types of facts directly in the prompting workflow without prompt guidance and demo examples; w/o FA indicates that after extracting the two types of facts from each case, the similarity is directly labeled without the guidance of positive and negative demos in the original FA progress.
Methods MF Kappa LF Kappa
Original 0.62 0.67
w/o ADM 0.45 0.58
w/o FE 0.21 0.30
w/o FA 0.26 0.38
w/o FE&FA 0.19 0.26
First, in a vertical comparison presented in Table 2, BERT trained only on the 2k synthetic dataset already shows significant improvements. Performance further improves substantially with the larger 20k dataset, indicating that our method can make effective relevance judgments. The improvement on the 20k synthetic data is shown across all baselines, demonstrating the general effectiveness of the synthetic supervision of our method. Notably, both BERT and Longformer, when trained on the 20k synthetic dataset, outperform the original SOTA model SAILER, highlighting the strong potential of our approach despite not relying on any additional complex post-processing techniques. Interestingly, despite doubling the data size, the performance on the 40k dataset was noticeably worse than on the 20k dataset. Two potential explanations are as follows: (1) a training dataset more aligned with the test set distribution yields greater benefits for the model; and (2) random sampling without filtering or selection introduces more unnecessary negatives and fewer positive examples, with excessive negatives adding limited value to the model. Then, in a horizontal comparison, we observed a phenomenon similar to that in Table 3: small LLMs fine-tuned on data augmented by a strong LLM using our workflow, when further used to annotate data within the same workflow for fine-tuning retrievers, achieve improvements that can approach or even surpass the performance of the strong LLM, With minimal expert supervision, our approach enables a large language model to acquire case analysis expertise, and subsequently transfers this ability to a smaller model via annotation-based knowledge distillation.
4.5. Ablation Study (RQ4)
@@
In this section, we aim to explore the impact of each stage in the proposed workflow on the final accuracy of the relevance judgments.
@@
According to Table LABEL:aba, removing Adaptive Demo-Matching (ADM) impacts the relevance judgments of Material Facts (MF) notably. This indicates that the extraction and relevance assessment of MF relies on more pertinent expert guidance. This may be due to the variable nature of MF. Meanwhile, Legal Facts (LF) remain unaffected due to their normative formats. This also highlights that the understanding of LF in the case benefits more from the Fact Extraction(FE) and Fact Annotation (FA) modules. When FE is removed, the Kappa scores for both types of facts significantly decrease, and the effect is similar when both FE and FA are removed. This demonstrates that the fact extraction step we designed is the most crucial part of the entire judgment process. Based on this observation, it can be inferred that LLM faces difficulty in accurately identifying key legal elements. Removing FA has little impact on LF annotation, but it does negatively affect MF. This indirectly confirms that extracting MF from case pairs and assessing their relevance is a challenging aspect of judging legal case relevance, and our method addresses this challenge through a series of carefully designed processes.
@@
Every stage within the workflow contributes significantly to enhancing the quality of legal relevance judgments. Both ADM and FA moderately influence the quality of relevance judgments, with the impact of ADM being more pronounced. ADM validates that relevant demonstrations indeed provide LLMs with a more robust basis for reasoning. On the other hand, FA introduces clearer constraint conditions for relevance judgments. Notably, the most pivotal process is FE, involving the most specialized knowledge Furthermore. It also suggests that our FE method can effectively leverage minimal expert instructions to inspire and guide LLMs in capturing subtle but vital nuances within professional data.
5. Limitation and additional experiments
Our study has two main limitations. First, we only evaluate on the case retrieval task, though it is highly relevant in real-world legal scenarios. Second, all experiments are conducted in Chinese, due to substantial differences between Chinese and foreign legal systems.
Despite this, our method generalizes well to other legal domains. We conducted additional experiments on CAIL2019333https://github.com/alumik/cail2019, which involves civil cases such as private lending, IP disputes, and maritime law. Unlike LeCaRD, each sample contains a query case and two candidates, requiring a binary relevance judgment.
Table 4. Results for CAIL2019 dataset
Model SOTA BERT BERT+20k
F-score 0.739 0.436 0.750
Using only 9 expert-designed demonstration sets, we generated pairwise annotations from unlabeled cases in the training set and selected 20k samples to fine-tune BERT. As shown in Table4, results compared with reported SOTA(Feng et al., 2024; Bi et al., 2022) show that our method remains effective in different legal domains.